OpenStudy (anonymous):

In June, Katherine bought 21 cans of food and 3 bags of litter for a total of $42. In July she spent $60 on 32 cans of food and 4 bags of litter. What is the cost of one bag of litter?

OpenStudy (anonymous):

Same thing as your previous question. What do you think the equations are?

OpenStudy (anonymous):

21x + 3y = 42 32x + 4y = 60 3x + y = 6 y = 6 - 3x 4y = 24 - 12x 32x + 24 - 12x = 60 21x = 60 -24 21x = 36 x = 36/21 you should be able to do the rest from here

OpenStudy (anonymous):

thankss